OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = x + 2. y = 4x - 1 y = 4x - 3 y = x - 3 y = x - 1

OpenStudy (anonymous):

if it parallels to a line, they have the same slope. so the slope of the new equation is also 1

OpenStudy (anonymous):

now you just need to find b in equation y=x+b when x=-4, y=-2

OpenStudy (anonymous):

so its y=4x-3?

OpenStudy (anonymous):

so in a equation, y=ax+b, a is always the slope, so if the slope is 1, a=1

OpenStudy (anonymous):

im so confused :/

OpenStudy (anonymous):